undeadknight26 (undeadknight26):

is this right? An airplane ascends at the rate of 40 feet per minute. The function below shows the height of the airplane from the ground f(t) after t minutes. f(t) = 40t + 520 Which statement best describes the y intercept of the graph of the function? It is equal to 520 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 560 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 40 and represents the height of the airplane above the ground, in feet, before it starts ascending.

undeadknight26 (undeadknight26):

It is equal to 520 and represents the height of the airplane above the ground, in feet, before it starts ascending.

OpenStudy (anonymous):

x=time (t) y=height f(x)

undeadknight26 (undeadknight26):

so 520 reps height then right?

undeadknight26 (undeadknight26):

SO it would be d? It is equal to 520 and represents the height of the airplane above the ground, in feet, before it starts ascending.

OpenStudy (anonymous):

yes that is the "y intercept" if you will of y=mx+b

undeadknight26 (undeadknight26):

so then it would be D right?

undeadknight26 (undeadknight26):

Since ti says that is the height before take off?

OpenStudy (anonymous):

It is equal to 520 and represents the height of the airplane above the ground, in feet, after 1 minute.

undeadknight26 (undeadknight26):

So the first one?

undeadknight26 (undeadknight26):

Did you know the answer to the other one i posted?

OpenStudy (anonymous):

mx+b m = slope, b = y intercept 40t + 520 40 = slope, 520 = y intercept

undeadknight26 (undeadknight26):

doesnt y = b also?

OpenStudy (anonymous):

link to it please?

undeadknight26 (undeadknight26):

The table below represents the distance a person runs as a function of time. Time (second) x Distance (meters) f (x) 10,60 13,78 16,96 19,114 The average rate of change of the function between x = 10 to x = 16 is __________ m/s and represents the person's speed.

undeadknight26 (undeadknight26):

I think it is -36

undeadknight26 (undeadknight26):

@ehuman

undeadknight26 (undeadknight26):

@HaleyT @ash2326 @helppls123 @phi @Preetha please help me...

undeadknight26 (undeadknight26):

Thanks for coming do you think this is the right answer?

undeadknight26 (undeadknight26):

@ehuman come back please!!!

undeadknight26 (undeadknight26):

@soulflower @austinL @amistre64 @Skyz @bloopman please help...

undeadknight26 (undeadknight26):

@hartnn

OpenStudy (anonymous):